You are an analyst employed by an airplane manufacturer that last year sold 40,000 ATR-72


aircrafts at $100,000 each. Your market research indicates that:


I. the price elasticity of demand for your aircrafts in −0.5. (or +0.5 in absolute value);


II. the income elasticity of demand for your aircrafts is +3.7; and


III. the cross price elasticity for your aircrafts with respect to the price of a comparable jet


manufactured by a competitor is +1.6.


A. Suppose that you expect a ceteris paribus decrease in average incomes of 10% this


year compared to last year. How many aircrafts do you estimate that your company will


sell this year? How will it impact total revenues? 6 mark

1
Expert's answer
2021-11-30T16:42:49-0500

"E(ID)=\frac{\Delta Quantity}{\Delta Income}"


"\frac{Quantity2 - Quantity2}{Quantity2+Quantity1} = -37"


"\frac{\Delta Quantity}{-10} = 3.7"

"\Delta Quantity2=-37"


"-37=\frac{Quantity2-Quantity2}{QQuantity2+QQuantity2}\\
=\frac{Quantity22-40000}{Quantity22+40000}"

"-38Quantity2=-1520 000"

"Quantity2=40000\\
Quanitity2=40000"